OpenStudy (anonymous):

Find the horizontal or oblique asymptote of f(x) = 2x^2+5x+6/x+3 a. y = -3 b. y = 2x - 1 c. y = 5 d. y = -2x + 5 @satellite73 @jim_thompson5910 @Directrix @SolomonZelman @iambatman @Callisto @wio

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

you get a quotient and a remainder ignore the remainder, the quotient is your "oblique" asymptote

OpenStudy (anonymous):

synthetic division is easiest in this case

OpenStudy (anonymous):

how do i do that? can you show me?

OpenStudy (anonymous):

but you can do long division too if you want

OpenStudy (anonymous):

list the coefficients they are 2 5 6

OpenStudy (anonymous):

can you show me please how to?

OpenStudy (anonymous):

then put a -3 on the side 2 5 6 -3 ___________

OpenStudy (anonymous):

bring down the 2 2 5 6 -3 ___________ 2

OpenStudy (anonymous):

then \(-3\times 2=-6\) put it under the 5 2 5 6 -3 -6 ___________ 2

OpenStudy (anonymous):

add and get \(5-6=-1\) 2 5 6 -3 -6 ___________ 2 -1

OpenStudy (anonymous):

now you can continue to get the remainder, but you do not need the remainder only the quotient you read off the coefficients from the bottom line

OpenStudy (anonymous):

can you continue? i dont know what to do after?

OpenStudy (anonymous):

divide a polynomial of degree 2 by a polynomial of degree 1, get a polynomial of degree 1 so answer is \(2x-\)

OpenStudy (anonymous):

oops \\[y=2x-1\] is what i meant to write

OpenStudy (anonymous):

2 5 6 -3 -6 ___________ 2 -1 y = 2x -1